OpenStudy (anonymous):

4sin(x)^2 + 1=8cos(x) Solve for x

OpenStudy (anonymous):

\[4(\sin x)^{2}+1=8cosx\]

OpenStudy (anonymous):

convert sin^2x to cos^2x first

OpenStudy (anonymous):

4(sin^2x) +1 =8cosx 4(1-cos^2x) +1 =8cosx 4cos^2x +8cosx-5=0

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

4cos^2x +10cosx-2cosx-5=0 2cosx(2cosx+5)-1(2cosx+5)=0 (2cosx-1)(2cosx+5)=0

OpenStudy (anonymous):

cosx <=1 hence cosx=1/2 x= 2n(pi)+- (pi/3) where n is any integer